OpenStudy (anonymous):

Find the length of Acr AB: if the angle measure is 65 degrees and the diameter is 14in

OpenStudy (anonymous):

The proportion of the arc to the circumference is the same as the proportion of the angle represented to the full 360 degrees. \[arc length / \pi*d = 65/360\] Rearrange for arc length....

OpenStudy (anonymous):

Arclength = Radius * Angle in Radians Arclength = 7 * 65(pi/180)